OpenStudy (anonymous):

c^2-2cd-80cd^2= (c+10d)(c-8d)??

jimthompson5910 (jim_thompson5910):

you mixed up your signs should be c^2-2cd-80cd^2= (c-10d)(c+8d)

OpenStudy (anonymous):

wait my signs are wrong

OpenStudy (anonymous):

lol yea i just noticed that lol thanks

jimthompson5910 (jim_thompson5910):

yep because -10+8 = -2 is what we want not 10+-8 = 2

OpenStudy (anonymous):

3x^2-20x-7...(3x+1)(x-7)? im a little confused on this one

OpenStudy (anonymous):

i figured it out.

jimthompson5910 (jim_thompson5910):

First multiply the first coefficient 3 by the last term -7 to get -21 So we need two numbers that BOTH: a) multiply to -21 AND b) add to -20 (the middle coefficient) It turns out that these two numbers are -21 and 1

jimthompson5910 (jim_thompson5910):

oh ok

jimthompson5910 (jim_thompson5910):

to check your work, you can FOIL out the factored answer and you should get the original expression back

OpenStudy (anonymous):

i thought the 3x^2 would throw me off

jimthompson5910 (jim_thompson5910):

it might, but if you step through it carefully, then it won't be so bad